OpenStudy (anonymous):

Need some help on this. Lol A company produces accessories for smart phones and tablets. The profit on each smart phone case is $5 and the profit on each tablet case is $3. The company made a profit of $900 on the cases last month. The equation 5x + 3y = 900 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. 1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

y = mx +c ( slope intercept form) eqn 5 x + 3y = 900 3y = 900 - 5x y= (900/3) - (5/3)x y = (-5/3)x + 300 this is the eqn in slope intercept form. here m is the slope therefore slope(m)= -5/3 hope it will help you.